# Using matrix-product states for time-series machine learning

### v1

@@ II.1 Encoding time series as product states
The MPS ansatz is a powerful tool for compressing functions in an exponentially large Hilbert space. To leverage the MPS formalism to represent the joint probability distribution of a time-series process, we must map the continuous-valued time-series amplitudes, $x_{t}$, to vectors in a finite-dimensional Hilbert space, to make the time-series data compatible with the MPS formalism. To achieve this, we employ $d$-dimensional nonlinear feature maps, which throughout the text we refer to as the encoding:
@@
which for a given choice of real or complex basis functions $b_{i}$, map a real valued time-series amplitude $x_{t}$ to a vector in $\mathbb{R}^{d}$ or $\mathbb{C}^{d}$. This mapping via the encoding $\phi_{t}(x_{t})$ is depicted schematically in Fig. 2(a). In general, the choice of basis functions $[b_{1},\ldots,b_{d}]_{t}$ can be time dependent, but here we consider a time-independent feature map. When applied to an entire time series of $T$ samples, the complete feature map $\Phi(\mathbf{x})$ is then given by the tensor product of local feature maps:
@@
To ensure the probabilistic interpretation of the wavefunction parameterized by the the matrix product state $W$ in Eq. (2) remains valid, the normalization condition:
@@
must be satisfied [15]. This condition will hold, provided that: (i) the MPS is unitary; and (ii) the feature map is orthonormal under the measure $\mu(x)$ [20]. While many valid choices of feature map exist, including a Fourier basis, Laguerre and Hermite polynomials, in the present work we choose either $b_{i}(x)=P_{i}(x)$ or $b_{i}(x)=CF_{i}(x)$, the $i$th orthonormal Legendre polynomial, or $i$th complex Fourier basis function, respectively. In particular, the Legendre polynomials are well suited to encoding time-series data, as they satisfy the orthonormality conditions on the compact interval $[-1,1]$, and can generalize to any physical dimension $d$. Higher physical dimensions enable more complexity within each site and, when coupled with larger bond dimension, more complex the inter-site MPS interactions. This comes at the cost of increased computational storage and runtime.
To encode real-world time-series data using the Legendre basis, we first apply a pre-processing step to map the original values $\mathbf{x}\in\mathbb{R}$ to a compact domain $\mathbf{x}\in[-1,1]$. When using the MPS for time-series classification, we apply a scaled outlier-robust sigmoid transformation [36], before transforming the dataset to map the dataset minimum to $-1$, and the dataset maximum to 1 (min–max normalization). For imputation, preserving the shape of the data is essential, and care must be taken to ensure minimal distortion to imputed time-series when transforming between the encoding and original data domain. For this reason, we avoid nonlinear transformations (e.g., sigmoid transform), which can amplify encoding-related errors, particularly when mapping from $[-1,1]\rightarrow\mathbb{R}$. Rather, we only apply a linear min-max normalization. Further details about the specific data transformations used here are in Appendix A.
@@ II.2 Training an MPS-based generative model
In the previous section, we laid the conceptual foundations for an MPS-based approach to time-series analysis. We motivated the MPS as being a powerful representation for encapsulating complex joint distributions underlying time series, and we described how we can represent time series in a form that is compatible with the MPS framework. In this section, we now outline the training procedure used to fit an MPS ansatz to a joint probability distribution given a finite training set of time series sampled from it. As we will show, the resulting MPS can be used to solve time-series machine learning problems, such as imputation of missing data, generation of new data, and classification of unseen data within a unified, interpretable, and tractable framework.
@@ II.2.1 Loss function for generative modeling
Given a dataset of $N$ time series, the goal of training is to learn an MPS model which closely approximates the joint probability distribution of the data, shown schematically in Fig. 2(b). Starting with a randomly initialized MPS, a common approach to training is to iteratively update the MPS entries by minimizing the Kullback–Leibler (KL) divergence [37, 20, 16, 38, 39], a measure of the discrepancy between the probability distribution $p(\mathbf{x})$ captured by the MPS, and the true distribution $q(\mathbf{x})$ that gave rise to the data:
@@
where the integral is over every possible configuration of $\mathbf{x}$. In practice, the true distribution $q$ is unknown, so it is common to minimize an averaged negative log-likelihood (NLL) loss function [20, 40, 21, 18],
@@
This is equivalent to minimizing the KL divergence (up to a constant). Once the MPS is trained, the joint probability density $p$ is extracted from the MPS using the Born rule, which now takes the form,
@@
To extend training to a classification setting, we treat each class of time series as originating from a separate joint probability distribution. Given $L$ classes of data, we need to learn $L$ probability distributions using $L$ separate MPSs. As in Stoudenmire and Schwab [15], an equivalent, yet more efficient approach is to attach an $L$-dimensional label index to a single site of an MPS:
@@
This allows us to approach generative modeling and classification in a unified manner, as the total loss $\mathcal{L}$ reduces to Eq. (7) in the single class ‘unsupervised’ case. To avoid confusion about ‘class dependence’ in the generative modeling sections, we will drop the label index $l$ from our notation when it is one dimensional.
@@ II.2.2 Sweeping optimization algorithm
In this section, we outline the local optimization algorithm we used to minimize $\mathcal{L}$, Eq. (10), on the training data. We use the same per-site DMRG-inspired sweeping algorithm in Stoudenmire and Schwab [15], but using a modified version of tangent-space gradient optimization (TSGO) [39] in the local tensor update step. Here, we present only the key algorithmic steps as a summary, and direct readers to Stoudenmire and Schwab [15] for a detailed treatment.
Starting with a randomly initialized MPS of $T$ sites, encoding dimension $d$, and uniform bond dimension $\chi_{\rm init}$.
@@
    Repeat steps 2-5 for each adjacent pair of remaining tensors, traversing from right-to-left, then from left-to-right, for a fixed number of iterations, or until convergence is achieved.
@@
As in Stoudenmire and Schwab [15], we define one sweep of the MPS to be a ‘round-trip’, i.e., one backward and one forward pass through all tensors in the MPS.
II.3 Imputation
@@
Many empirical time series contain segments of time for which values are missing (e.g., due to sensor drop-out) or contaminated (e.g., with artifacts). This complicates resulting analyses and, for many algorithms (such as classification or dimension-reduction), requires imputation, i.e., that the missing data be inferred (‘filled in’) from values that were observed [42, 43]. For example, in some astrophysical star surveys, periods of sensor downtime is inevitable as a result of essential operational procedures [44], leaving ‘gaps’ in the recorded light curves of stars during which no observations are made. Existing approaches to imputing missing values in time series range from filling them with a simple statistic of observed values (e.g., the mean of the full time series [45]), linearly interpolating across the missing data, to inferring missing values based on fitted linear models (e.g., autoregressive integrated moving average (ARIMA)), similarity-based methods (e.g., $k$-nearest neighbor imputation from a training set of data (K-NNI) [46]), and machine learning algorithms (e.g., generative adversarial networks (GANs) [47], variational auto-encoders (VAEs) [48]). In this work, we present a generative approach to time-series imputation by leveraging the data distribution encoded by the MPS $W$ (as in Eq. 2), summarized schematically in Fig. 3 using Penrose graphical notation [49]. Our imputation approach, shown schematically in Fig. 2(c), consists of two main steps: (i) conditioning the MPS on known values; and (ii) inferring missing values from the conditioned MPS. Here, we summarize the key algorithmic steps and refer the reader to Appendix D for full details.
@@
Figure 3: An MPS-based algorithm for time-series imputation. Here we consider an illustrative example of an imputation problem involving a six-site MPS, represented graphically with Penrose notation [49], where the time-series values $x_{2},x_{4},x_{6}$ are observed, and we would like to impute the unobserved values $x_{1},x_{3},x_{5}$. (a) Two copies of the trained MPS (one conjugate-transposed, indicated by the dagger, $\dagger$) with open physical indices encode the joint distribution over all possible states, given by $\rho_{1,2,\dots,6}$ (Eq. (14)). (b) The MPS is projected into a subspace where the states $s_{2}$, $s_{4}$, $s_{6}$, corresponding to each of the known time-series values, have been measured. The updated MPS now encodes the joint distribution over the remaining states $s_{1}$, $s_{3}$, $s_{5}$, conditional upon having measured $s_{2},s_{4},s_{6}$. (c) The single site conditional reduced density matrix $\rho$ (as in Eq. (15)) is obtained by tracing over all remaining sites. By evaluating the probability density function $\textrm{pdf}_{i}(x)=\phi_{i}^{\dagger}(x)\rho_{i}\phi_{i}(x)$ for $x$ in the encoding domain, we estimate the value for $x_{1}$ and its uncertainty $\Delta x_{1}$ using the median (see Eq. (17)) and weighted median absolute deviation (WMAD), respectively. (d) The MPS is projected onto the imputed state $s_{1}=\phi_{1}(x_{1})$ (Eq.(15)), yielding an updated MPS which encodes the joint distribution over $s_{3},s_{5}$. (e, f, g) We repeat estimating the next missing value using the median of the pdf (conditioned on all previously known and imputed values), then projecting the MPS onto the corresponding states, until all missing values are recovered.
@@
Let the ordered sequence $\mathbf{x}^{*}=x_{1},x_{2},\ldots,x_{T}$ represent a time series with known observations $\{\mathbf{x}^{o}:o\in O\}$, and missing values $\{\mathbf{x}^{m}:m\in M\}$, where $M:=\{1,2,\ldots,T\}\setminus O$. The reduced density matrix for the unknown states $\rho_{M}$ can be obtained by first projecting the observed states onto the MPS:
@@
where $\widetilde{W}^{\dagger}$ is the conjugate transpose of $\widetilde{W}$. Throughout the text, we refer to the process of projecting the MPS onto a known value as either a measurement or conditioning a joint distribution on a known value. The scale factors $\sqrt{P(x_{i})}$ ensure that the MPS remains correctly normalized. Technical details of the implementation are described fully in Appendix D.
An important feature of MPS-based imputation is that each missing value is imputed one at a time, using both known observations and already imputed values that were previously missing. Therefore, the outcome of an imputation depends on the order in which it is performed. Time series are ordered in time, so we elect to perform imputation sequentially from earliest to latest using the ‘chain rule’ of probability, as described in [50]. While previous approaches typically sample a random state from the single-site conditional distributions (e.g., via inverse transform sampling [20]), here we use a deterministic approach wherein we select the state that corresponds to the median of the conditional distribution at each imputation site. Our choice to use the median is motivated by the fact that discretizing continuous time-series values using a finite number of basis functions $d$ artificially broadens the probability distribution. Sampling directly from this broadened distribution can artificially lead to outlying values being chosen which degrades the imputation performance. As detailed in Appendix B, selecting the median of the distribution effectively mitigates this broadening effect and provides a single best estimate for each imputation site. In what follows, we present our MPS-based algorithm for time-series imputation, recapitulated in Fig. 3. Starting at the earliest missing site, $i$, in $M$:
@@
    Infer $x^{*}_{i}$ using the median of the conditional distribution
@@
    Project the MPS onto the selected state $s_{i}$ by contracting the MPS with $\phi_{i}^{\dagger}(x_{i}^{*})$, as in Eq. (13):
@@
    Proceed to the next unknown time point and repeat steps 1–4 using the lower-rank MPS $\widetilde{W}^{\prime}$ until all missing values $\mathbf{x}^{m}$ are recovered.
II.4 Classification
Having established the theoretical foundations for training an MPS to learn a generative distribution underlying an observed time-series dataset, here we show how our framework can be naturally extended to the supervised setting to address the problem of time-series classification – the second key application of our approach. Specifically, for a dataset of time-series instances, each associated with a distinct class label $l$ from a finite set of classes $L$, the primary objective is to learn a model that can accurately predict the class label $l\in L$ of new, unseen time series by leveraging the learned generative distributions of the classes.
As stated in Sec. II.2, classification is achieved by training $L$ distinct MPSs. For each class, indexed by $l$, we train a separate MPS $W$ on the subset of time-series instances belonging to $l$. To classify an unlabeled time series $\mathbf{x}$, we then compute its (non-normalized) probability density $p(\mathbf{x})$ under each of the $L$ MPSs, and select the label $l$ of the MPS that assigns the highest relative probability to the instance, as illustrated schematically in Fig. 2(d).

### v2

@@ II.1 Encoding time series as product states
The MPS ansatz is a powerful tool for compressing functions in an exponentially large space. To represent the joint probability distribution underlying a time-series dataset using MPS, the continuous-valued time-series amplitudes, $x_{t}$, must first be mapped to vectors in a finite-dimensional Hilbert space. To achieve this mapping, we employ $d$-dimensional nonlinear feature maps, which throughout the text we also refer to as the ‘encoding’:
@@
which for a given choice of real or complex basis functions $b_{i}$, map a real valued time-series amplitude $x_{t}$ to a vector in $\mathbb{R}^{d}$ or $\mathbb{C}^{d}$, respectively. This mapping via the encoding $\phi_{t}(x_{t})$ is depicted schematically in Fig. 2(a). In general, the choice of basis functions $[b_{1},\ldots,b_{d}]_{t}$ can be time dependent, but here we consider a time-independent feature map. When applied to an entire time-series instance of $T$ samples, the complete feature map $\Phi(\mathbf{x})$ is then given by the tensor product of local feature maps:
@@
To ensure the probabilistic interpretation of the wavefunction parameterized by the the Matrix-Product State (MPS) $W$ in Eq. (2) remains valid, the normalization condition:
@@
must be satisfied [15]. This condition will hold, provided that: (i) the MPS is unitary; and (ii) the feature map is orthonormal under the measure $\mu(x)$ [20]. While many valid choices of feature map exist, including the Fourier basis, Laguerre and Hermite polynomials, in the present work we choose $b_{i}(x)=P_{i}(x)$, the $i$-th orthonormal Legendre polynomial. The Legendre polynomials are well suited to encoding time-series data, as they satisfy the orthonormality conditions on the compact interval $[-1,1]$, and can generalize to any physical dimension $d$. Higher physical dimension $d$ gives rise to larger tensors which, coupled with larger bond dimension, allows the MPS to express more complex inter-site interactions. Such increased model complexity, however, comes at the cost of additional computational storage and runtime.
To encode real-valued time-series amplitudes using the Legendre basis, we first apply a pre-processing step to map the original values in the data domain $\mathbf{x}\in\mathbb{R}$ to a compact encoding domain $\mathbf{x}\in[-1,1]$. When using the MPS for time-series classification in Sec. III.3 we apply a scaled outlier-robust sigmoid transformation [39], before transforming the dataset to map the dataset minimum to $-1$, and the dataset maximum to 1 (via min–max normalization). For imputation, preserving the shape of the data is essential, and thus care must be taken to ensure minimal distortion to imputed time series. For this reason, we avoid nonlinear transformations (e.g., sigmoid transform), which can amplify encoding-related errors, particularly when mapping from the encoding domain to the original data domain i.e., $[-1,1]\rightarrow\mathbb{R}$. Instead, for time-series imputation, we apply a linear min–max normalization as the only pre-processing step. Further details about the specific data transformations we use are in Appendix A.
@@ II.2 Training an MPS-based generative model
In the previous section, we described how continuous-valued time-series instances can be converted into a discrete representation suitable for MPS-based machine learning. We now detail the procedure used to train an MPS in order to model the joint distribution defined by a finite set of time-series instances sampled from it. As we will show, the resulting MPS can then be used to solve diverse time-series machine learning problems, such as imputation of missing data, classification of unseen data, and generation of new data, within a general framework.
@@ II.2.1 Loss function for generative modeling
In order to approximate the joint distribution $p(\mathbf{x})$ of the data, an MPS model is trained on a finite number of encoded time-series instances sampled from it, shown schematically in Fig. 2(b). Starting with a randomly initialized MPS, the model is trained by iteratively updating its tensor entries to minimize the Kullback–Leibler (KL) divergence [40, 20, 16, 41, 42], a measure of the discrepancy between the probability distribution $p(\mathbf{x})$ captured by the MPS, and the true distribution $q(\mathbf{x})$ from which the training data are sampled:
@@
where the integral is over every possible configuration of $\mathbf{x}$. In practice, the true distribution $q$ is unknown, so it is common to minimize an averaged negative log-likelihood (NLL) loss function over the set of available training time-series instances $N$ [20, 43, 21, 18],
@@
This is equivalent to minimizing the KL divergence (up to a constant). Once the MPS is trained, the joint probability density $p(x)$ is extracted from the MPS using the Born rule, which now takes the form,
@@
To extend training to classification, we treat each class of time series as originating from a separate joint probability distribution. In other words, given $L$ classes of data, classification requires learning $L$ probability distributions with $L$ unique MPS models. As in Stoudenmire and Schwab [15], an equivalent, yet more efficient approach is to attach an $L$-dimensional label index to a single site of an MPS:
@@
The addition of a label index $l$ allows us to approach generative modeling and classification in a unified manner, as the total loss $\mathcal{L}$ reduces to Eq. (7) in the single class ‘unsupervised’ case. To avoid confusion about class dependence in the generative modeling sections, we will omit the label index $l$ from our notation when it is one dimensional (i.e., single class).
@@ II.2.2 Sweeping optimization algorithm
Using the loss function $\mathcal{L}$ in Eq. (10) to quantify the MPS approximation of the data joint distribution, we now outline the local optimization algorithm we used to minimize its value on the training data. We use the same DMRG-inspired sweeping algorithm in Stoudenmire and Schwab [15], but with a modified version of tangent-space gradient optimization (TSGO) [42] for the local tensor update step. Here, we present only the key algorithmic steps as a summary, and direct readers to Stoudenmire and Schwab [15] for a detailed treatment.
Starting with an MPS of $T$ sites, with each site being represented by a tensor of random entries, fixed encoding dimension $d$, and uniform initial bond dimension $\chi_{\rm init}$.
@@
    Repeat steps 2-5 for each adjacent pair of remaining tensors, traversing from right-to-left, then from left-to-right, for a fixed number of iterations, or until convergence of the loss function $\mathcal{L}$ is achieved.
@@
As in Stoudenmire and Schwab [15], we define one sweep of the MPS to be a ‘round-trip’, i.e., one backward and one forward optimization pass through all tensors in the MPS.
II.3 Time-series imputation algorithm
@@
Many empirical time series contain segments of time for which values are missing (e.g., due to sensor drop-out) or contaminated (e.g., with artifacts). This complicates analyses and, for many algorithms (such as classification or dimension-reduction), requires imputation, i.e., that the missing data be inferred (i.e., ‘filled in’) from values that were observed [45, 46]. For example, in some astrophysical star surveys, periods of sensor downtime is unavoidable due to operational procedures [47], leaving ‘gaps’ in the recorded light curves of stars during which no observations are made. Existing approaches to imputing missing values in time series range from filling them with a simple statistic of observed values (e.g., the mean of the full time series [48]), linearly interpolating across the missing data, to inferring missing values based on fitted linear models (e.g., autoregressive integrated moving average (ARIMA)), similarity-based methods (e.g., $k$-nearest neighbor imputation from a training set of data (K-NNI) [49]), and machine learning algorithms (e.g., generative adversarial networks (GANs) [50], variational auto-encoders (VAEs) [51]). In this work, we present a generative approach to time-series imputation by leveraging the data distribution modeled by the MPS $W$ (as in Eq. 2), summarized schematically in Fig. 3 using Penrose graphical notation [52]. Our imputation approach, shown schematically in Fig. 2(c), consists of two main steps: (i) conditioning the MPS on observed values, and (ii) using the resulting conditioned MPS to infer missing values. Here, we summarize the key algorithmic steps and refer the reader to Appendix D for full details.
@@
Figure 3: An MPS-based algorithm for time-series imputation. Here we consider an illustrative example of an imputation problem involving a six-site MPS, represented graphically with Penrose notation [52], where the time-series values $x_{2},x_{4},x_{6}$ are observed, and we would like to impute the unobserved values $x_{1},x_{3},x_{5}$. (a) Two copies of the trained MPS (one conjugate-transposed, indicated by the dagger, $\dagger$) with open physical indices encode the joint distribution over all possible states, given by $\rho_{1,2,\dots,6}$ (Eq. (14)). (b) The MPS is projected into a subspace where the states $s_{2}$, $s_{4}$, $s_{6}$, corresponding to each of the known time-series values, have been measured. The updated MPS now encodes the joint distribution over the remaining states $s_{1}$, $s_{3}$, $s_{5}$, conditional upon having measured $s_{2},s_{4},s_{6}$. (c) The single-site conditional reduced density matrix $\rho$ (as in Eq. (15)) is obtained by tracing over all remaining unmeasured sites. By evaluating the probability density function $\textrm{pdf}_{i}(x)=\phi_{i}^{\dagger}(x)\rho_{i}\phi_{i}(x)$ for $x$ in the encoding domain, we estimate the first unobserved value $x_{1}$ and its uncertainty $\Delta x_{1}$ using the median (see Eq. (17)) and weighted median absolute deviation (WMAD) of the probability density function, respectively. (d) The MPS is projected onto the estimated state $s_{1}=\phi_{1}(x_{1})$ (Eq.(15)), yielding an updated MPS which encodes the joint distribution over remaining unmeasured states $s_{3},s_{5}$. (e, f, g) We repeat the process of estimating the next missing value (i.e., $x_{3}$) using the median of the corresponding pdf (conditioned on all previously known and imputed values), then projecting the MPS onto the estimated state (i.e., $s_{3}=\phi_{3}(x_{3})$), until all remaining unobserved values are recovered.
@@
Let the ordered sequence $\mathbf{x}^{*}=x_{1},x_{2},\ldots,x_{T}$ represent a time series with observed values $\{\mathbf{x}^{o}:o\in O\}$, and missing values $\{\mathbf{x}^{m}:m\in M\}$, where $M:=\{1,2,\ldots,T\}\setminus O$. The reduced density matrix for the unknown states $\rho_{M}$ can be obtained by first projecting the observed states onto the MPS:
@@
where $\widetilde{W}^{\dagger}$ is the conjugate transpose of $\widetilde{W}$. Throughout the text, we refer to the process of projecting the MPS onto an observed value as either a measurement or conditioning a joint distribution on an observed value. The scale factor $\sqrt{P(x_{i})}$ ensure that the MPS remains correctly normalized throughout. Technical details of the implementation are described fully in Appendix D.
An important feature of MPS-based imputation is that each missing value is imputed one at a time, using both observed values and imputed values that were previously missing. Therefore, the outcome of an imputation depends on the order in which it is performed. As time series are indexed in time order, we choose to perform imputation sequentially from earliest to latest using the ‘chain rule’ of probability, as described in [53]. While previous approaches typically sample a random state from the single-site conditional distributions (e.g., via inverse transform sampling [20]), we use a deterministic approach wherein we select the state that corresponds to the median of the conditional distribution at each imputation site, as in Fig. 3(c). Our choice to use the median is motivated by the observation that discretizing continuous time-series values using a finite number of basis functions $d$ induces an artificial broadening of the MPS-approximated probability distribution. Sampling directly from this broadened distribution can produce values that are unlikely under the true data distribution, leading to degraded imputation performance. As detailed in Appendix B, selecting the median of the distribution effectively mitigates this broadening effect and provides a robust best estimate for each imputed value. In what follows, we present our MPS-based algorithm for time-series imputation, shown schematically in Fig. 3. Starting at the earliest missing site, $i$, in $M$:
@@
    Infer the unobserved value $x^{*}_{i}$ using the median of the conditional distribution as:
@@
    Project the MPS onto the selected state $s_{i}$ corresponding to the inferred value $x^{*}_{i}$ by contracting the MPS with $s_{i}=\phi_{i}^{\dagger}(x_{i}^{*})$, as in Eq. (13):
@@
    Proceed to the next unobserved value’s MPS site and repeat steps 1–4 using the lower-rank MPS $\widetilde{W}^{\prime}$ until all unobserved values $\mathbf{x}^{m}$ are recovered.
II.4 Time-series classification algorithm
Building on the formalism in Sec. II.2, we now describe how the MPSTime algorithm can be extended to the problem of time-series classification. Specifically, for a dataset of time-series instances, each associated with a distinct class label $l$ from a finite set of classes $L$, we train a model that can accurately predict the class label $l\in L$ of new, unseen time-series instances using the learned generative distributions of the classes.
As described in Sec. II.2, classification is achieved by training $L$ distinct MPSs. For each class indexed by $l$, we train a separate MPS $W$ on the subset of instances belonging to $l$. To classify an unseen time-series instance $\mathbf{x}$, we then compute its (non-normalized) probability density $p(\mathbf{x})$ under each of the $L$ MPSs, and select the label $l$ of the MPS that assigns the highest relative probability to the instance, as illustrated schematically in Fig. 2(d).